### v2

# Active Label Correction for Building LLM-based Modular AI Systems

Abstract
Large Language Models (LLMs) have been used to build modular AI systems such as HuggingGPT, Microsoft Bing Chat, and more. To improve such systems after deployment using the data collected from human interactions, each module can be replaced by a fine-tuned model but the annotations received from LLMs are low quality. We propose that active label correction can be used to improve the data quality by only examining a fraction of the dataset. In this paper, we analyze the noise in datasets annotated by ChatGPT and study denoising it with human feedback. Our results show that active label correction can lead to oracle performance with feedback on fewer examples than the number of noisy examples in the dataset across three different NLP tasks.
Active Label Correction for Building LLM-based Modular AI Systems
Karan Taneja Georgia Institute of Technology karan.taneja@cc.gatech.edu                        Ashok Goel Georgia Institute of Technology ashok.goel@cc.gatech.edu
1 Introduction
Large language models (LLMs) such as OpenAI’s ChatGPT OpenAI (2022), Google’s Bard based on LaMDA Thoppilan et al. (2022), and Meta’s LLaMA Touvron et al. (2023) are powerful zero or few-shot learners as they can generalize to a wide range of tasks simply through task explanations or demonstrations without any model fine-tuning Qin et al. (2023). This has led to the development of many modular AI agents that leverage these LLMs through API services to decompose complex tasks and automate workflows. Some popular examples include HuggingGPT Shen et al. (2023), AutoGPT Significant Gravitas (2023), and Bing Chat. Further, many enterprises are building their own conversational AIs based on LLMs as zero-shot learners as well as response-generation models Tellez (2023). These agents rely on several discriminative and generative sub-tasks to achieve the goal specified by the user. For example, one of HuggingGPT’s sub-tasks is to perform model selection from HuggingFace Model Hub based on model descriptions. While producing the final natural language output is a generative task, many sub-tasks like model selection are discriminative.
AI agents based on LLMs work reasonably well in practice, but their robustness is still in question as they are limited by the quality of zero-shot learning of the LLMs that they employ Shen et al. (2023). An error in one module can trigger a cascade effect, propagating the error throughout the modular system. Further, LLMs acting as zero-shot learners cannot be easily fine-tuned for the end tasks in every AI agent for two reasons: (i) a lack of training data, and (ii) the computational cost required to train and maintain an online service for multiple fine-tuned models can be expensive and also has a large carbon footprint. AI developers also cannot simply rely on future improvements in the quality of underlying LLMs as the performance on certain end-tasks can also worsen Chen et al. (2023) likely due to fine-tuning that focuses on the safety and conversationality of LLMs.
In this paper, we discuss a pipeline based on active label correction Rebbapragada et al. (2012) for improving modular AI systems that rely on LLMs as zero-shot learners and we evaluate its efficacy for three NLP tasks. We propose utilizing the data traces containing user inputs and LLM outputs from the initial deployment to learn a smaller substitute model for each module. We employ active label correction which methodically uses human expertise by seeking annotations for examples that are most likely misannotated by LLM. This process encourages leveraging the powerful zero-shot learning capabilities of the LLMs to rapidly deploy modular AI agents and improve their quality as they accumulate data from the real world. We focus on discriminative tasks in our experiments for this paper but the proposed method also extends to generative tasks in principle (see Appendix A).
Theoretically, the optimization problem at our hand is to select the optimal subset of examples that, when corrected by human annotators, will jointly minimize the test error and the cost of correction. Formally, let $\theta$ be the parameters of a machine learning (ML) model $M$ trained using dataset $\mathcal{D}$. $E(\theta)$ be the error of the model over the test set, and $C(\mathcal{D}_{m})$ be the cost of annotating $\mathcal{D}_{m}$ (for example, $C(\mathcal{D}_{m})=|\mathcal{D}_{m}|$). From a dataset $\mathcal{D}$ with noisy annotations, we wish to find the subset $\mathcal{D}_{m}\subseteq\mathcal{D}$ of examples for which we obtain human annotations or corrections to form $\mathcal{D}_{(m)}$. The optimization problem can be written as
$\min_{\mathcal{D}_{m}\subseteq\mathcal{D},\theta}\quad$ $E(\theta)+\eta\cdot C(\mathcal{D}_{m})$
s.t. $\theta=M(\mathcal{D}\setminus\mathcal{D}_{m}\cup\mathcal{D}_{(m)})$ (1)
where $\eta$ is a trade-off factor. This is similar to the machine teaching problem Zhu (2015) for which a closed-form solution does not exist in the general case, and especially in LLMs where model parameters are uninterpretable and the input domain of text is extremely complex. We use active label correction as an iterative greedy approach by alternately (i) seeking corrections for most likely misannotated examples based on predictions from a trained model, and (ii) training the model to learn from the updated annotations.
Our paper makes three main contributions: (i) we propose a pipeline based on active label correction to continually improve modular AI agents while minimizing annotation costs, (ii) we compare performance-cost curves of different ALC methods to study their efficacy on datasets annotated using ChatGPT as a zero-shot learner, and (iii) we present experiments for three NLP tasks that are common to modular AI agents including natural language inference, named entity recognition, and intent classification.
2 Related Work
LLM-based Modular AI Agents: ChatGPT OpenAI (2022) and LaMDA Thoppilan et al. (2022) are conversational LLMs trained using large amounts of text and fine-tuned with human feedback. ChatGPT has been shown to perform well on many diverse NLP tasks using prompts Qin et al. (2023) though it doesn’t perform at par with fine-tuned models. LLaMa Touvron et al. (2023) is an open-source LLM trained for zero-shot learning and surpasses GPT-3 on many benchmarks. HuggingGPT Shen et al. (2023) is a modular AI agent that utilizes ChatGPT to solve complex tasks like manipulating objects in image, video, text, and audio domains using a large collection of models from HuggingFace [^1] https://huggingface.co/. AutoGPT Significant Gravitas (2023) is another example of an AI agent based on ChatGPT and GPT-4 to solve tasks by planning, goal setting, task prioritization, and task execution. These agents depend on numerous calls to LLMs behind API services and are expensive to run. We propose a pipeline to use data traces from the deployment of AI agents and limited human supervision to improve their performance and to make them cost-effective by using smaller models. We use ChatGPT as our zero-shot learner as it is easily available as an API service and it performs well at following instructions. We found BARD based on LaMDA Thoppilan et al. (2022) to be challenging to use as it consistently ignores the instructions related to the output formatting.
Figure 1: Proposed process for improving LLM-based modular AI agents. A noisy annotated dataset obtained from a zero-shot or few-shot learner-based module is fed to a misannotation predictor to get flagged examples. The flagged examples are corrected by a human annotator and used to update the dataset. This process is performed iteratively until a stopping condition is met and the final dataset is used to train a substitute model.
Semi-supervised Learning and Active Learning: In semi-supervised learning, training data consists of a set of labeled examples and a typically much larger set of unlabeled examples Hady and Schwenker (2013). Weakly supervised learning more generally assumes that annotations are partial, noisy, or provided at a higher level of abstraction Zhou (2018). Self-training is a common method used to train a teacher model on the supervised subset to predict soft pseudo-labels for the unsupervised subset. Then, the supervised subset and pseudo-labeled subset are used to train a student model Karamanolakis et al. (2021). This can be done iteratively and end-to-end to improve pseudo-labels and model parameters Yi and Wu (2019); Tanaka et al. (2018); Yao et al. (2019). On the other hand, active learning is a strategy to select the most informative examples, typically by constructing them or choosing from a pool of unlabeled examples, and querying an oracle for supervision Settles (2009). In our work, we begin with a fully annotated but noisy dataset and evaluate the performance gain and cost reduction associated with iteratively seeking human expertise for improving data and model quality.
Active Label Correction (ALC): ALC, introduced by Rebbapragada et al. (2012), uses the self-training process to clean a noisy training set by iteratively predicting most likely mislabeled examples using a trained model and receiving human feedback to correct them. They find that simple disagreement measures based on the output probabilities work well for identifying mislabeled examples. Robust ALC Kremer et al. (2018) assumes that noise is conditioned on class labels and uses a latent model for the true label. Dual ALC or DALC Li et al. (2022) additionally uses model predictions with high confidence for automatic correction in their work. These works have primarily focused on image classification problems in vision and experimented with synthetic noise. We explore the use of ALC and DALC for correcting noise in ChatGPT’s text annotations. Our proposed pipeline is aimed at leveraging noisy data to improve LLM-based modular AI systems.
3 Methods
Modular AI agents similar to HuggingGPT Shen et al. (2023) rely on the zero-shot learning power of LLMs to process the input, perform sub-tasks, and generate output. These agents can accumulate large amounts of data including inputs, intermediate outputs of each module, and final outputs through their interactions with users. To evaluate a module that relies on LLMs as a zero-shot learner, one can analyze the inputs and outputs amassed by the module. Figure 1 outlines our pipeline. We propose collecting input-output pairs for each module to create a noisy annotated dataset, and improving data quality using human feedback with the goal of training a replacement model for the zero-shot learner. The dataset formed by input-output pairs is noisy because the output predictions from zero-shot learners have lower quality than that of a human-annotated dataset. Using ALC Rebbapragada et al. (2012), we iteratively predict and flag the examples that are most likely to be noisy or misannotated, and seek correction from human annotators on the flagged examples. This iterative greedy approach to solving Equation 1 is motivated by (i) the goal of minimizing noisy training examples which is a natural proxy for test set error, and (ii) using minimum human feedback by effectively predicting misannotated examples.
3.1 Misannotation Prediction
The task of misannotation prediction (MP) involves predicting the examples in a dataset that are most likely to contain noise based on the entire noisy dataset. We wish to train a model that learns patterns from the noisy dataset and flags examples that do not conform with these patterns.
Given a noisy dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ where $(x_{i},y_{i})$ are pairs of inputs and outputs, we train a model $p_{\theta}$ to predict $y^{*}=\arg\max_{y}p_{\theta}(y|x)$. We then obtain predictions on the training dataset itself and store the probabilities $p_{\theta}(y_{i}|x_{i})$ for each example. Similar to Rebbapragada et al. (2012), we suppose that an example is misannotated with probability $m(x_{i},y_{i})=1-p_{\theta}(y_{i}|x_{i})$. In other words, if the probability of example $(x_{i},y_{i})$ is low under the trained model, it is very likely to be misannotated ($m\approx 1$) and vice versa. We refer to this method as ALC.
Intuitively, when the output for a training example is inconsistent with the annotation in the dataset, we can argue that the remaining examples suggest evidence against the given annotation and, therefore, it is likely to be misannotated. Since each example has a minimal impact on the model for a large enough sample size, $p_{\theta}(y_{i}|x_{i})$ approximates as the probability of the model trained with one example removed from the dataset. We study this impact of data size in the experiments and results section.
We also experiment with Random Label Correction or RLC as a baseline where the examples are effectively assigned a random $m$-value i.e. $m(x_{i},y_{i})=\text{Random}(0,1)$.
While ALC employs human supervision for correction, the model predictions can also be a source of correct annotations for a large proportion of examples. Dual Active Label Correction (DALC), introduced by Li et al. (2022), leverages these predictions by using both the model predictions and human feedback to update the dataset. We adapt their method and compare it to ALC by using model predictions to change labels before determining misannotated examples. We accept output label $y^{*}$ as true annotation for an example if $p_{\theta}(y^{*}|x)>\delta$ where $\delta$ is a constant threshold. We refer to this automatic update of annotations based on model predictions as auto-flipping as it requires no human input. For the remaining examples, we use ALC to find the most likely misannotated examples. We shall refer to this method as DALC.
Flagged Examples: After computing $m(x_{i},y_{i})$ $\forall$ $(x_{i},y_{i})\in\mathcal{D}$, we sort $(x_{i},y_{i})\in\mathcal{D}$ from highest to lowest value of $m(x_{i},y_{i})$. We flag $M\in(0,1)$ fraction of examples in $\mathcal{D}$ with the highest $m(.)$ values to be misannotated. The fraction $M$ can be changed to control the number of flagged examples. A low value of $M$ can lead to a low recall as many misannotated examples will be missed in the flagged set while a high $M$ can mean lower precision because many correctly annotated examples may be included in the flagged set. In DALC, we do not flag examples that are auto-flipped.
3.2 Iterative Data Correction
The flagged examples obtained from the misannotation predictor are fed to a human annotator for correction. Once corrected by a human expert, these examples can be used to update the dataset. We can perform another iteration of MP including the model training on the updated dataset to obtain a new set of flagged examples. This process can be repeated to iteratively improve the quality of the data until a stopping criterion is met. Stopping conditions will be discussed in the results section. For DALC, in each iteration, we revert the previously auto-flipped examples and auto-flip a new set examples based on new model predictions. Further, an example corrected by human annotation cannot be auto-flipped or flagged in subsequent iterations.
With each iteration, new information obtained from the human annotator is accommodated into the dataset which is used to train the next model for MP. The new model predictions reveal new inconsistencies with the noisy annotations and lead to the next set of flagged examples. As the count of misannotated examples reduces in the noisy dataset with each iteration, we can expect the MP performance to drop. We discuss this phenomenon in the results section. One way to control the MP performance is to control $M$, the fraction of flagged examples. Ideally, one can train a new model after correcting every single example but this extreme case is expensive and infeasible due to the long wait time as MP runs after every example is verified. But we still may want to adjust the value $M$ based on a target precision value. Therefore, in practice, the MP precision can be measured for the latest batch of examples and $M$ can be decreased if precision is low or vice versa. Increasing precision by lowering $M$ comes at the cost of low recall and running more iterations. In our experiments, we use a fixed value of $M$ i.e. a fixed number of examples are flagged in each iteration.
4 Experiments and Results
4.1 Performance on Evaluation Tasks
We evaluate the ALC-based pipeline on three different tasks: (1) ATIS (Airline Travel Information System) Intent Classification Hemphill et al. (1990), (2) CoNLL 2003 Named Entity Recognition or NER Tjong et al. (2003) and (3) Question-answering Natural Language Inference or QNLI Wang et al. (2018). The three tasks are common to many conversational AI agents Lin et al. (2023). Intent classification is used for the downstream processing of incoming requests, NER is used for slot-filling (such as train booking) to fill various parts of the input before a database query, and NLI or textual entailment is used in conversational AI agents to confirm the validity of the output in answering the user query. All three datasets are publicly available and span different levels of complexity and sizes as shown in Table 1. Furthermore, the ground truth annotations allow us to compare our pipeline to the oracle performance.
For each task, we created a prompt using the OpenAI Playground and then used the OpenAI API to get outputs for both the train and test split of each dataset using ‘gpt-3.5-turbo’ model with zero temperature i.e. no randomness in output generation. For fine-tuning, we used ‘distilbert-base-uncased’, ‘roberta-base’ and ‘albert-large-v2’ pre-trained transformer models from HuggingFace for ATIS, CoNLL and QNLI tasks respectively based on their public leaderboards [^2] https://paperswithcode.com/ and availability on HuggingFace. The results for each task are shown in the last column FT in Table 1. Model training details are provided in Appendix B.
Task Metrics Train Eval FT
ATIS Size 4,952 878 878
Acc. 0.702 0.756 0.982
F1 Sco. 0.789 0.832 0.982
CoNLL Size 14,041 3,250 3,250
#NERs 23,499 5,942 5,942
Precision 0.576 0.614 0.958
Recall 0.757 0.765 0.961
F1-Sco. 0.654 0.682 0.959
QNLI Size 104,743 5,463 5,463
Acc. 0.849 0.845 0.920
Table 1: Performance of ChatGPT on train and test set of three tasks: (1) ATIS Intent Classification, (2) CoNLL 2003 Name Entity Recognition, and (3) QNLI Natural Language Inference along with performance of fine-tuned models in FT column.
ATIS dataset consists of 17 unique intents (‘airport’, ‘aircraft’, ‘distance’, etc.) assigned to user queries. A description of all 17 intents was provided to ChatGPT for classification. 40 examples with multiple labels were removed from the dataset to reduce the task to single-label classification. For the CoNLL dataset, an example of entity recognition output was provided in the prompt to describe the format and four types of entities i.e. person, location, organization, and miscellaneous. QNLI dataset consists of pairs of questions and sentences based on the Stanford Question Answering Dataset Wang et al. (2018). We provide the inputs in the prompt and ask if the given passage can be used to answer the question. Prompts and output parsing details are provided in Appendix C.
The performance of ChatGPT on the three tasks are documented in Table 1. Overall, the results agree with Qin et al. (2023) in terms of the difference in performance between a fine-tuned model and zero-shot learning using GPT-3. We observed a large improvement in F1 Score on CoNLL 2003, 68.2% compared to F1 Score of 53.7% in Qin et al. (2023), possibly because of our different prompt instruction which explicitly asks for the classification of each word instead of extracting a list of named entities. We note that ChatGPT has 10-30% lower performance than fine-tuned models across the three tasks. While zero-shot learners such as ChatGPT generalize well to new tasks without training, this result suggests that there is a large performance gap that needs to be bridged when bootstrapping modular AI agents with ChatGPT. Therefore, it is important to develop a method to improve each module and the overall system over time after deploying the LLM-based AI systems.
Noise Accuracy C. Imbalance
None 98.2% -
Random 97.2% 0.1697
Label-cond. 93.5% 0.0485
Input-cond. 88.6% 0.1033
ChatGPT 84.0% 0.0949
Table 2: Comparison of different noise types, including no noise (None), Random, Label-conditional, Input-conditional, and ChatGPT annotations, on model Accuracy and Class Imbalance for ATIS dataset.
4.2 Noise Characteristics
We compare different types of noise models to understand the noise induced by ChatGPT while annotating the ATIS dataset. We chose the ATIS dataset for this experiment as it has the highest number of classes which can lead to more pronounced noise patterns. We add different types of synthetic noise to the original dataset with the same proportion of misannotations as ChatGPT. Table 2 shows the accuracy of models trained with different noisy versions of ATIS datasets and the class imbalance measured by the KL divergence of label distribution from the original label distribution. Details about noise models are provided in Appendix E.
We find that ChatGPT adds the most detrimental noise while having a similar class imbalance as other noise models. We believe that ChatGPT provides very reasonable false annotations that are inherently hard to detect or ignore which leads to poor test performance. It is also interesting to note that a model trained on ChatGPT-annotated training data performs much better than ChatGPT on the eval dataset (84.0% versus 75.6%). This is likely due to the input data distribution learned by the model leading to better input representations as compared to a generic pre-trained model. This motivates the fine-tuning of task-specific models for applications in modular AI systems as well as the use of active label correction to improve data quality which can lead to further improvement in the performance of modular AI agents.
Figure 2: A 2D projection of ATIS text embeddings for a subset of 7 classes. ChatGPT annotations are indicated by colors while large dots indicate errors.
In Figure 2, we show 2D t-SNE projections Maaten and Hinton (2008) of Contriever text embeddings Izacard et al. (2022) from ATIS dataset. We visualize examples labeled with 7 classes shown in the legend and larger dots indicate the examples misannotated by ChatGPT. We observe that most misannotated examples lie near intersections or between class clusters. This indicates that noise induced by ChatGPT is indeed complex and depends on the text, its true label, and other possible labels. This also conforms with Table 2 where we find random and label-conditional noise to be non-representative models of ChatGPT annotations.
4.3 Misannotation Prediction
In Figure 3, we compare RLC, ALC, and DALC described earlier for MP on all three datasets annotated using ChatGPT. We use $\delta=0.75,0.98,0.90$ for ATIS, CoNLL and QLNI to keep the proportion of auto-flipped examples between 3% to 6%. We note from Table 1 that 29.8% of the ATIS and 15.1% of the QNLI train dataset were misannotated by ChatGPT. For CoNLL, 57.4% of examples contain errors in token classes though only 9.7% of tokens in the dataset are misclassified.
Figure 3: Misannotation Prediction (MP) precision and recall for ATIS, CoNLL, and QNLI with change in $M$.
We observe that precision goes down for ALC and DALC as $M$ increases for each task. This is because the examples with the highest probability of being misannotated based on $m(.)$ values indeed have a higher rate of being misannotated. In other words, starting from the examples with the highest $m(.)$ values, the human annotator will see a decrease in the MP quality over time as they start verifying examples with lower $m(.)$ values. We also see an increase in recall as $M$ increases which is expected because more misannotated examples are flagged at higher values of $M$. We also note that DALC has consistently lower precision and recall than ALC. The auto-flipped examples in DALC typically have high $m(.)$ as $p_{\theta}(y\neq y^{*}|x)$ is low but these examples cannot be flagged. ALC, on the other hand, flags the examples with the highest $m(.)$ values leading to higher precision and recall. For RLC, the precision is constant and low while recall increases perfectly linearly with $M$ because examples are flagged at random.
For the human annotator, examples flagged by DALC are more likely to be correct as compared to those flagged by ALC. It provides a more balanced set of flagged examples for the correction task which may also avoid the problem of over-reliance of human annotators on model predictions. Further, the overall quality of the data can be higher with DALC because of two sources of true labels viz. auto-flipping and human corrections. Given a specific method, we wish to maximize precision to correct a higher proportion of misannotated examples in each iteration. Therefore, we use low values of $M$ in each step in iterative data correction.
4.4 Effect of Data Size on MP
We wish to understand how the quantity of data collected by modules within an AI system will impact MP performance. Towards this, we vary the amount of data used to train the model for MP. We use 100%, 50%, 25%, 12%, 6%, and 3% of the QNLI dataset, the largest of the three with 105k examples (3% $\approx$ 3.1k examples), to evaluate the impact of data size on the MP performance. The results are shown in Figure 4 and 5.
Figure 4: MP precision versus flagged examples ($M$) for fractions (3%-100%) of QNLI dataset.
Figure 5: MP precision versus the percentage of the QNLI dataset used for different values of $M$.
We observe a decrease in the MP precision as the size of the dataset decreases in Figure 4. We mentioned earlier that $p_{\theta}(y|x)$ can be approximated as the output probability of a model trained with the leave-one-out dataset when the sample size is large enough. With a decrease in data size, each example plays a more significant role in the learning process. We confirm this as the drop in precision from 6% to 3% is much bigger than drop from 100% to 50%.
In Figure 4, we plot the precision as a function of data size with a fixed $M$. For each $M$, the precision increases as the data size increases but with diminishing returns. In practice, the observed precision can suggest the data size increase required to improve MP performance.
Figure 6: Improvement in model performance (Accuracy/F1-Score) with iterations of simulated human verification using RLC, ALC, and DALC along with MP Precision and the Oracle performance.
4.5 Iterative Data Correction
The quality of training data improves after each iteration of label correction and the iterative process can be continued until we obtain a high-quality dataset. We simulate correction, without actual human annotators, by using the ground truth known from the original dataset to update the annotations of the flagged data. We fix $M=0.025$ (2.5%) for ATIS and QNLI tasks, but $M=0.050$ (5.0%) for CoNLL because $>50\%$ examples contain noisy tokens. Further, for CoNLL task, we report the F1 score on token-level classification as a standard practice. However, we predict misannotations on the sentence level since the human annotators verify complete sentences containing many tokens. Therefore, we calculate misannotation precision by treating every token as flagged if a sentence is predicted to be misannotated. While this leads to seemingly low precision, it is more useful as a trend compared to sentence-level precision which is consistently high because most examples for CoNLL contain misannotated tokens.
We train models with the updated datasets and plot performance i.e. Accuracy or F1 (using $\times$ markers) after each iteration in Figure 6. We also plot the MP precision for each iteration (using $\bullet$ markers). We say that a model has achieved close-to-oracle performance if the accuracy or F1 score is within a 1% range of the model trained on the original dataset (Oracle). The horizontal green bar in each plot shows this 1% range. For the ATIS task, DALC achieves close-to-oracle performance in 11 iterations. With $M=2.5\%$, this means that the human annotator examines $2.5\%\times 11=27.5\%$ of the dataset which is less than the $29.8\%$ of data misannotated by ChatGPT. ALC takes 2 more iterations i.e. another $2.5\%\times 2=5.0\%$ of data needs correction. For CoNLL, we again observe that DALC requires 11 iterations to achieve close-to-oracle performance. Since $57.4\%$ of CoNLL annotations by ChatGPT contain tokens with incorrect classes, it is reasonable for human annotators to observe $5.0\%\times 11=55\%~{}(<57.4\%)$ of the data. ALC requires another 2 iterations or $10\%$ data-equivalent. For QNLI, we observe close-to-oracle performance after 6 iterations using DALC i.e. correcting $2.5\%\times 6=15\%~{}(<15.1\%)$ of the dataset. ALC requires another iteration or $2.5\%$ data-equivalent. For RLC, the MP precision remains constant as examples are randomly selected for correction. It leads to a much slower improvement in model performance compared to other methods.
We also observe a gradual decrease in MP precision for ALC and DALC due to a decrease in the number of misannotated examples after each iteration. As human annotators examine the flagged examples in practice, we can calculate the MP precision after each iteration and use it as a stopping criterion or as a signal to decrease $M$. Additionally, we hold out a random human-annotated test set to evaluate the model performance after each iteration. If the MP precision is no longer in a target range or the model performance saturates, one can stop performing iterations until more data becomes available from the deployed AI agent.
Overall, we found that DALC can achieve close-to-oracle performance by correcting only a fraction of the dataset. We find that this fraction is less than the fraction of misannotated examples for all three ChatGPT-annotated datasets.
5 Conclusion
Modular AI agents can leverage LLMs as zero-shot or few-shot learners for rapid deployment without training but there is a large performance gap when compared to fine-tuned models. The proposed pipeline utilizes ALC to improve modular AI agents by leveraging LLM-annotated data and improving its quality to train a replacement model for each module. Compared to synthetic noise models used in previous applications of ALC to the vision domain, we found the noise characteristics of ChatGPT annotated text data to be much more complex. We achieve the same performance as ground truth data using human feedback on a fraction of the dataset. Furthermore, we found this fraction to be less than the fraction of misannotated data for three unique tasks spanning different levels of complexity and data sizes.
6 Limitations
(i) Most publicly available datasets have small text lengths while real applications may involve much longer (conversational) texts. Deploying an AI agent and collecting real datasets is outside the scope of this paper. (ii) Most commercial AI agents allow users to provide feedback through a thumbs-up/down button and this feedback may be accommodated as a signal for misannotation prediction in modular AI agents. We leave this as an open problem for future work. (iii) In real applications, outliers or completely irrelevant examples should be removed from the training dataset, and exploring methods for identifying such examples in LLM-annotated data is another problem that should be addressed in the future work.
Appendix A Extension to Generative Models
To generalize our approach to generative tasks, we can use the conditional probability of generated text given a text input.
$P(y|x)=P(y_{1},\ldots,y_{n}|x)=P(y_{1:n}|x)$
For decoder models, sampling is performed auto-regressively by assuming $P(y_{1},\ldots,y_{n}|x)=P(y_{1}|x)\cdot P(y_{2}|y_{1},x)\ldots P(y_{n}|y_{1:n-1},x)$. After model training with a dataset of input and output texts, the probability of misannotation can be calculated as usual.
$m(x,y_{1:n})=1-P(y_{1:n}|x)$
Generative tasks are far more complex than discriminative tasks. One can expect that the amount of data required to train a model for misannotation prediction for generative tasks will be much more than for discriminative tasks. In future work, we will examine the data size required to effectively perform annotation prediction for different generative tasks.
Appendix B Training Details
The training details for models for each task are provided in Table 3.
ATIS CoNLL QNLI
Model Name distilbert-base-uncased roberta-base albert-large-v2
Trainable Parameters 67M 124M 17M
Training examples 4,952 14,041 104,743
Epochs 9 6 3
Batch Size 16 16 16
Optimizer AdamW AdamW AdamW
Learning Rate 1e-4 (constant) 2e-5 (constant) 1e-5 (constant)
Warmup Steps 155 439 0
Table 3: Training details for ATIS, CoNLL and QNLI tasks.
Appendix C ChatGPT Prompts for Evaluation Tasks
C.1 ATIS
ATIS (Airline Travel Information System) Intent Classification dataset Hemphill et al. (1990) is labeled with 17 unique intents with one or more intents assigned to user queries. There are only 40 examples with multiple labels. We removed them from the dataset leading to a single-label classification task. We provided ChatGPT with the system message below to instruct it for classification.
‘role’: ‘system’ ‘content’: ‘Your task is to classify the query into possible classes.’
Along with the system message, we provide the user message to describe the problem, the input text, and the output format.
‘role’: ‘user’ ‘content’: ‘Classes: [Class 1]: [class description], [Class 17]: [class description], Query: [input text] Answer exactly one of [class names separated by commas].’
A few examples of classes and their descriptions are as follows:
  1. 1.
    ‘aircraft’: ‘This intent is associated with queries that seek information specifically about an aircraft, such as its features, type, or specifications.’
  2. 2.
    ‘flight’: ‘This intent covers general queries related to flights, including information about schedules, availability, or any other flight-related inquiries.’
  3. 3.
    ‘flight time’: ‘Queries with this intent typically seek information about the duration or specific timing of a flight.’
  4. 4.
    ‘cheapest’: ‘This intent indicates queries that focus on finding the most economical or low-cost options, such as the cheapest flights or fares.’
  5. 5.
    ‘ground fare’: ‘Queries with this intent are related to ground fares or transportation costs, such as taxi fares or shuttle services.’
To retrieve the intent from the output, we iterate over the intents and use the first found intent in the output text.
C.2 CoNLL
To extract named entities from examples in CoNLL 2003 dataset Tjong et al. (2003), we provided ChatGPT with one example of entity recognition output to describe the output pattern. First, we used the user message below to provide the types of entities that need to be extracted.
‘role’: ‘system’ ‘content’: ‘Your task is to extract named entities of type Person, Location, Organization and Miscellaneous.’
Next, we use the user message below to provide an example and the input text from which entities need to be extracted. We use the same NER example below for every input.
‘role’: ‘user’ ‘content’: ‘Example input: “Mercury was born Farrokh Bulsara in Stone Town in the British protectorate of Zanzibar (now part of Tanzania) on 5 September 1946 ." Example output: “[Mercury](Person) was born [Farrokh Bulsara](Person) in [Stone Town](Location) in the [British](Miscellaneous) protectorate of [Zanzibar](Location) (now part of [Tanzania](Location)) on 5 September 1946 ." Input: [input text]’
To parse the output, we use regex to extract entity types and align words in the input and output to get classification for each token.
C.3 QNLI
QNLI (Question-Answering Natural Language Inference) dataset Wang et al. (2018) has pairs of questions and passages based on the Stanford Question Answering Dataset and the label 00 if the passage answers the question and $1$ otherwise. We use the system message below to prime ChatGPT with the task.
‘role’: ‘system’ ‘content’: ‘Your task is to solve a textual entailment task.’
Next, we use the following user message to give the input question and passage and prompt ChatGPT to say ‘YES’ if the passage can answer the question and ‘NO’ otherwise.
‘role’: ‘user’ ‘content’: ‘Question:[input question] Passage: [input passage] Can the given passage be used to answer the question? Answer exactly YES or NO (capitalized).’
Finally, we label the example with 00 if ‘YES’ is present in the output and $1$ otherwise.
Appendix D Misannotation Prediction for CoNLL
For CoNLL, each sentence is provided with a list of tokens and token-level classification. But we perform misannotation prediction (MP) on the sentence level since the human annotators evaluate the complete sentence and assign named entity types based on the context. Therefore, we use combine the probabilities of token-level misannotations into a sentence-level misannotation probability. Let $j$th token of sentence $x_{i}$ be $w_{j}$ with named entity label $y_{i}^{w_{j}}$. Let there be $L_{i}$ tokens in the sentence $x_{i}$.
An example is correctly annotated if all the tokens in the sentence are correctly labeled. Therefore, the probability of correct sentence-level annotation is the product of token-level probabilites.
$p_{\theta}(y_{i}|x_{i})=\prod_{1\leq j\leq L_{i}}p_{\theta}(y_{i}^{w_{j}}|x_{i})$
But leads to a lower probability for longer sentences. To normalize for length, we use the following equation:
$p_{\theta}(y_{i}|x_{i})=\Big{(}\prod_{1\leq j\leq L_{i}}p_{\theta}(y_{i}^{w_{j}}|x_{i})\Big{)}^{1/L_{i}}$
which is the geometric mean of token-level probabilities.
Appendix E Noise Models
For random noise, we randomly select examples for the noising process. For each selected example, we change the label to any other label with equal probability. In other words, we assume that each example is equally likely to be mislabeled and each label is equally likely to be replaced by another label.
For label-conditional noise, we first train a classification head on top of the BERT model (‘bert-base-uncased’ on HuggingFace) and use its probability outputs on training examples to calculate label transition probabilities. For each label, we calculate the transition probabilities by averaging the output distribution of examples with that label. After this, we randomly select examples for noising and change their labels based on the transition probabilities. This process assumes that each example is equally likely to be noisy but noisy label is conditionally dependent on the true label. See Li et al. (2022) (Section 3.1) for an introduction to label transition matrix as a representation for label-conditional noise model.
For input-conditional noise, we use the same BERT + fine-tuned classification head model to calculate the output probabilities of each example. We iterate over the dataset to sample new labels for each example based on its output probability distribution. We iterate over random permutations of the dataset multiple times but only sample new labels for unchanged examples. This is repeated until the required number of noisy examples is achieved. In this process, an example is likely to be mislabeled with probability proportional to $1-p_{\theta}(y_{i}|x_{i})$. Further, the new label is sampled with probability proportional to $p_{\theta}(y\neq y_{i}|x_{i})$.